OpenStudy (anonymous):

Can someone please help me?? A lottery winner must decide between two methods of payment. Choice one is to receive a lump sum of $100,000,000. Choice two is to begin with one cent on day 1, four cents on day 2, 16 cents on day three, and so on until the end of 18 days. Which choice should the lottery winner select? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

Do you know how to use exponents, e.g. 2^3?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

What's the equation for the amount of money that the person gets on Day X?

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

Each day that goes by he gets 4 times as much money as the last one. So that looks like an exponential function, rigth?

OpenStudy (anonymous):

What exponential function gives you f(0) = 1, f(1) = 4, f(2) = 16?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

It has to have an X in it to represent the day number

OpenStudy (anonymous):

2^x

OpenStudy (anonymous):

if x=1, what's 2^x? What about if x=2?

OpenStudy (anonymous):

if x=1then2^x is 2, if x=2 then 2^x = 4

OpenStudy (anonymous):

ok, that doesn't seem right then, right? We wanted when x=1, it's 1, and when x=2, it's 4

OpenStudy (anonymous):

so try again, can you think of a function that goes up by a factor of 4 each time?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4? You need an X for it to be a function